The first step to solving the equation 3 (x + 7) = 12x – 51 is to
combine like terms.
add the opposite to both sides.
distribute.
none of the above.
The second step to solving the equation 2x – 7 – (x – 5) = 0 is to
combine like terms.
add the opposite to both sides.
distribute.
divide both sides.

Found 2 solutions by ewatrrr, jim_thompson5910:
Answer by ewatrrr(24785)

 
Hi,
3 (x + 7) = 12x – 51
3x + 21 = 12x - 51 | Step I: Distribute

2x – 7 – (x – 5) = 0
2x - 7 - x + 5 = 0 | Step I: Distribute
x - 2 = 0 | Step II: Combine Like Terms

Answer by jim_thompson5910(35256)

The first step to solving the equation 3 (x + 7) = 12x – 51 is to
combine like terms.
add the opposite to both sides.
distribute.
none of the above.

You start off by distributing to get 3x + 21 = 12x - 51

From there, you need to move all the x terms to one side and get everything else to the other. Then you simplify. Afterwards, you can completely isolate x.
-------------------------------------------------------
The second step to solving the equation 2x – 7 – (x – 5) = 0 is to
combine like terms.
add the opposite to both sides.
distribute.
divide both sides.

The first step is to distribute (like in the first problem before this). This will give you

2x – 7 – x + 5 = 0

Then you combine like terms to get

x - 2 = 0

I'm sure you see what to do from here (if not, let me know)


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------